# Gaining Wisdom from Setbacks[Uncaptioned image]: Aligning Large Language Models via Mistake Analysis

### v3

@@ Appendix B More on Attack Defending
In Table 6, we present the experimental results for all baseline models using mistakes derived from ChatGLM. These findings are in alignment with our previous observations.
@@ Appendix D More Analysis on Equation (2)
which simplifies to $p(\bm{T}|\bm{Y},\bm{X})\propto p(\bm{Y}|\bm{X},\bm{T})$ under the assumption that $p(\bm{Y}|\bm{X})$ remains relatively stable during the fine-tuning process. To maintain the model’s capability to produce a response given an instruction (i.e., preserving $p(\bm{Y}|\bm{X})$), we combine original SFT data, both helpful ($D_{\text{helpful}}$) and harmless ($D_{\text{harmless}}$), with mistake analysis data during fine-tuning, rather than solely relying on the mistake analysis data. Thus, Eqn. (2) actually aims to explore how the additional mistake analysis data, along with the original SFT instruction-response pairs, can enhance alignment performance.
@@ Appendix E More Qualitative Comparison
We provide more examples of mistake analysis generated by Alpaca-7B and ChatGLM-6B in Fig. 10-13(a), with Fig. 10 and Fig. LABEL:fig:analysis_data_origin_chatGLM for harmful responses of the original training corpus, and Fig. 11 and Fig. LABEL:fig:analysis_data_chatGLM_chatGLM for the model-generated ones. We observe both Alpaca-7B and ChatGLM-6B are proficient at identifying the harmful responses with comprehensive explanation, and thus, incorporating the mistake analysis during SFT would enhance the generative capacities of LLMs with their innate discriminating abilities.

### v5

@@ Appendix B More on Attack Defending
In Table 6, we present the experimental results for all baseline models using mistakes derived from ChatGLM. These findings are in alignment with our previous observations in Table 1.
@@ Appendix D More Analysis on Equation (2)
which simplifies to $p(\bm{T}|\bm{Y},\bm{X})\propto p(\bm{Y}|\bm{X},\bm{T})$ under the assumption that $p(\bm{Y}|\bm{X})$ remains relatively stable during the fine-tuning process. To maintain the model’s capability to produce a response given an instruction (i.e., preserving $p(\bm{Y}|\bm{X})$), we combine original SFT data, both helpful ($D_{\text{helpful}}$) and harmless ($D_{\text{harmless}}$), with mistake analysis data during fine-tuning, rather than solely relying on the mistake analysis data. Thus, Eqn. (2) actually aims to explore how the additional mistake analysis data, along with the original SFT instruction-response pairs, can enhance the alignment performance.
@@ Appendix E More Qualitative Comparison
We provide more examples of the mistake analyses generated by Alpaca-7B and ChatGLM-6B in Fig. 10-13(a), with Fig. 10 and Fig. LABEL:fig:analysis_data_origin_chatGLM for harmful responses of the original training corpus, and Fig. 11 and Fig. LABEL:fig:analysis_data_chatGLM_chatGLM for the model-generated ones. We observe both Alpaca-7B and ChatGLM-6B are proficient at identifying the harmful responses with comprehensive explanation, and thus, incorporating the mistake analysis during SFT would enhance the generative capacities of LLMs with their innate discriminating abilities.